OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!! Two sets of quadratic expressions are shown below in various forms: Line 1: x^2 − x − 6 (x + 2)(x − 3) (x + 0.5)^2 − 6.25 Line 2: x^2 − x − 2 (x + 1)(x − 2) (x − 0.5)^2 − 2.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (anonymous):

@AkashdeepDeb @ganeshie8 @mathmale @sidsiddhartha

OpenStudy (akashdeepdeb):

Just simply multiply them and check! Here is the hint: \[(a+b)^2 = a^2 + b^2 + 2ab\] \[(a-b)^2 = a^2 + b^2 - 2ab\] \[(x+a)(x+b) = x(x+b) + a(x+b) = x^2 + bx + ax + ab = x^2 + (a+b)x + ab\] Try using these to see if the forms are the same.

OpenStudy (anonymous):

I think it's just the second line.

OpenStudy (anonymous):

Ok, I just multiplied, is it line 2 only? cuz that was the only 1 i got all equivalents

OpenStudy (akashdeepdeb):

Correct! :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

thanks @Shaunticlair 2! :) id give u a medal 2 if i could :)